OpenStudy (anonymous):

What is the inverse of the given relation? y = 3x + 9 A. y = 1/3x + 3 B. y = 3x - 3 C. y = 3x + 3 D. y = 1/3x - 3

OpenStudy (anonymous):

its a.

OpenStudy (johnweldon1993):

so your inverse...is found by switching your x's and y's, then solving for y again you have y = 3x + 9 so x = 3y + 9 -9 -9 x - 9 = 3y ---------- 3 3 3 y = 1/3x-3 so it is actually D

OpenStudy (anonymous):

poo...johns right.-. i messed up!

OpenStudy (johnweldon1993):

haha...no problem @Chooch146 :)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

http://www.purplemath.com/modules/invrsfcn3.htm